OpenStudy (anonymous):

if three envelopes are selected at random without replacement, determine the probability that they all are red envelopes ... 7 envelopes in all 3 are red, 2 are blue,1 green, 1 yelllow

OpenStudy (anonymous):

Incomplete question... how many envelopes were in the box, and how many were red or other colours

OpenStudy (anonymous):

there are 7 envelopes 3 red 2 blue 1 green 1 yellow

OpenStudy (anonymous):

@fortune_paul

OpenStudy (anonymous):

The definition of probability is : possible outcomes/total outcome so we have 3 red envelopes out of 7 possible envelopes, therefore the probability of getting the red envelope = 3/7

OpenStudy (anonymous):

@dlewis5

OpenStudy (anonymous):

See i thought that at first then i looked at the back of the book the answer was 1/ 35

OpenStudy (anonymous):

@fortune_paul

OpenStudy (anonymous):

@dlewis5 okay, the probability of picking a red the first time is : 3/7 second time : 2/6 (note that the number of red has decreased and the number of total enveloped decreased) third time : 1/5 multiply them all out = 3/7 * 2/6 * 1/5 = 1/35